OpenStudy (anonymous):

(3x^2a-4y^az^3a)^2 simplified?

OpenStudy (kewlgeek555):

Okay, first you ahve to combine like terms.

OpenStudy (kewlgeek555):

Hold on.

OpenStudy (kewlgeek555):

Do you know how to combine like terms?

OpenStudy (anonymous):

not so sure? its algbra 2.

OpenStudy (kewlgeek555):

Oh - no wonder. I am like trying to figure this out and the exponents are confusing me. Its that I am in Algebra I. I know how to do it with out exponents. Hold on.

OpenStudy (kewlgeek555):

Is this what you are asking?:\[(3x ^{2}a-4y ^{a}z ^{3}a)^{2}\]

OpenStudy (kewlgeek555):

If so, \[9a ^{2}x ^{4}-24a ^{2}x ^{2}y ^{a}2^{3}+16a ^{2}y ^{2a}5^{6}\]No, that sounds wrong. I an sorry. I don't really know. Hopefully @PFEH.1999 @campbell_st @Firejay5 @savannaxx_ can help you, I am not sure. Sorry.

OpenStudy (anonymous):

did @kewlgeek555 rewrite the problem right?

OpenStudy (campbell_st):

you are basically expanding a perfect square \[(a -b)^2 = a^2 -2ab + b^2\] in your question \[a = 2x^{2a}\] and \[b = 4y^az^{3a}\] you will need to use the index law for power of a power \[(x^a)^b = x^{a \times b}\] the power operates on every part of the term as well as the law for multiplying the same base \[x^a \times x^b = x^{a + b}\] hope this helps...